OpenStudy (anonymous):

what is the domain f(x,y)= ((1-x^2-5y^4))

OpenStudy (anonymous):

are there supposed to be any square roots? last time in your problem you had a bunch of parenthesis, but they were really square roots.

OpenStudy (anonymous):

Ok :) i do the square :) wait me

OpenStudy (anonymous):

\[\sqrt{1-x^2-5y^4}\]

OpenStudy (anonymous):

Same idea as last time. We dont want the inside of that square root to be negative. so we want: \[1-x^2-5y^4 \geq 0 \iff 1\geq x^2+5y^4\]from there I dont really see where to go with it, let me think about it for a sec.

OpenStudy (anonymous):

please I'd appreciate it a lot

OpenStudy (anonymous):

i wait :) dont worry

OpenStudy (anonymous):

i guess maybe I could solve for y...that would give us: \[x^2+5y^4 \leq 1 \iff 5y^4\leq 1-x^2\iff y^4 \leq \frac{1-x^2}{5}\]\[\iff y \leq \sqrt[4]{\frac{1-x^2}{5}}\]

OpenStudy (anonymous):

why you change the sense > to <

OpenStudy (anonymous):

i undersand :)

OpenStudy (anonymous):

iam graphicing :)wait me

OpenStudy (anonymous):

do you know how graphic' :(